Paul is 11 years younger than Wendy. In 10 years, Wendy will be twice as old as Paul. How old is Paul now?
Answer by Maths68(1474)

Let
Present age of Paul = p years old
Present age of Wendy = w years old

Paul is 11 years younger than Wendy.
w=p+11............(1)
In 10 years,
Paul will be = p+10 years old
Wendy will be = w+10 years old
Then
Wendy will be twice as old as Paul
w+10=2(p+10)
w+10=2p+20
w=2p+20-10
w=2p+10................(2)
Put the value of w from (1) to (2)
p+11=2p+10
p-2p=10-11
-p=-1
p=1
Put the value of p in (1)
w=p+11............(1)
w=1+11
w=12

Present age of Paul = p = 1 year old
Present age of Wendy = w = 12 years old